OpenStudy (anonymous):

i need help solving 7*(sq rt of 24) - 3*(sq rt of 6)

OpenStudy (anonymous):

use a calculator?

jhonyy9 (jhonyy9):

7sqrt24 -3sqrt6 = 7sqrt(6*4) -3sqrt6 =7*2sqrt6 -3sqrt6 = 14sqrt6 -3sqrt6 = =11sqrt6

OpenStudy (anonymous):

\[7\sqrt{6*4}- 3\sqrt{6} \]

OpenStudy (anonymous):

\[14\sqrt{6} -3\sqrt{6}\] \[11\sqrt{6}\]

OpenStudy (anonymous):

https://www.wba2.aplusanywhere.com/R85Content/media/pictures/alg12_nbg/b3426t1.gif can u help solve this one

OpenStudy (anonymous):

Distrubute the \[3\sqrt{2}\]

OpenStudy (anonymous):

then it would be \[3\sqrt{4}+12\sqrt{12}\] \[6+12\sqrt{4*3}\] \[6+24\sqrt{3} \]